(1 point) Find y as a function of t if y - 8y + 12y = 0, y(0) = 3, y(1) = 9. yt) = Remark: The initial conditions involve vaI AM REALLY STRUGGLING ON THIS PROBLEM PLEASE HELP ME CORRECT AND NEAT WORK IS MUCH APPRECIATED THANKS
Answer #1

G(+) = ? * Y - By +12y = 0 , 900)-3 , 9(1)= 9, Find roots of characteristic equation : 72-88 +12 = 0 →! 72-67-2r +32 = 0 8(